• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name nanda

    Log entry time 10:44:32 on September12,2005

    Entry number 154494

    keyword=compton setup

    Compton polarimeter is now set up to take data.

    Shiftcrew should take 2 hr runs while monitoring counting rates. Make sure background rates are below 100 and signal+bkg is above 300 Hz/uA for both electrons and photons.